A boat accelerates at a rate of 3.5 m/s2 in the direction of north of west. Find components.
1
Expert's answer
2020-12-06T17:26:47-0500
"a_x=a\cos{45}=3.5\cos{45}=2.47\frac{m}{s^2}\\a_y=a\sin{45}=3.5\sin{45}=2.47\frac{m}{s^2}"